OpenStudy (anonymous):

2^2x + 2^(x + 2) - 32 = 0 how would you solve for x?

myininaya (myininaya):

\[2^{2x}+2^22^x-32=0\] 2^2 is 4 then let u=2^x you will notice all you have to is solve a quadratic

myininaya (myininaya):

Do you understand what I'm asking you to do @swift_13 ?

OpenStudy (anonymous):

why is it 2^2*2^x

myininaya (myininaya):

\[2^{x+2}=2^x \cdot 2^2 \text{ By law of exponents}\]

myininaya (myininaya):